OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry of the quadratic equation y = 2(x - 4)^2 + 7? x = 7 x = -4 x = 4 x = -7

OpenStudy (mertsj):

Remember the axis goes through the vertex and is a vertical line so; x=4

OpenStudy (anonymous):

y=a(x-h)^2+k axis of symmetry is x=h the axis of symmetry is 4

OpenStudy (anonymous):

the equation for axis of symmetry is x = -b/2a

OpenStudy (anonymous):

in the standard form y = ax^2 + bx + c

OpenStudy (anonymous):

yours is in vertex form

OpenStudy (anonymous):

so it makes it easier

hero (hero):

I awarded study and mridrik medals for posting good general solutions.

OpenStudy (anonymous):

thanks